# Controlled bit-flip of period-doubling and discrete time crystalline states in open systems

### v1

@@ III.3 Beyond parametric oscillator models
Figure 6: (a) Switching probability of the ODM for different particle number and for the parameters $\lambda_{0}=1.1\lambda_{c}$, $\kappa=1.0\omega$, $\omega_{d}=0.8\omega$, and $A=0.55$. (b)-(d) Half-winding number distribution of the ODM for (b) $N=10^{4}$, (c) $N=10^{5}$, (d) $N=10^{6}$. The same set of parameters in (a) is used to construct the histograms in (b)-(e).
We present in Fig. 6(a) the bit-flip probability of the DTC states in the $\lambda_{0}>\lambda_{c}$ regime. Notice that $P_{s}$ has a more complex behavior, with its most striking feature being the discontinuity at $T_{\delta,c}\approx 0.8T_{d}$ that persists even for small $N$. This is in contrast to the behavior observed so far with the thermal DPP and DTC states in the NP in Figs. 4(a) and 4(b), respectively. As we further increase $T_{\delta}$, however, $P_{s}$ drops to a lower value. In particular, when we neglect the temporal fluctuations and only consider the initial noise, $P_{s}\approx 0.5$ within the range of $T_{\delta}/T_{d}\in\left[1.2,1.5\right]$ and $T_{\delta}/T_{d}\in\left[1.7,2.0\right)$, while it goes to zero for $T_{\delta}/T_{d}\in\left(1.5,1.7\right)$. Unlike the discontinuous transition in $T_{\delta}=0.8T_{d}$ however, this feature does not persist for small $N$, and instead becomes smoothened out similar to the crossovers observed in Figs. 4(a) and 4(b).
From the distribution of $w$ shown in Figs. 6(b)-6(d), we see that at $T_{\delta,c}=0.8T_{d}$, the distribution jumps from $w=0$ to $w=+1$ and then splits into $w=+1$ and $w=-1$ as we further increase $T_{\delta}$. This splitting of the $P_{w}$ persists even for large values of $N$, which, again, is in contrast to the thermal DPP and the DTC states for $\lambda_{0}<\lambda_{c}$, where the distribution only becomes localized at one value of $w$. This splitting persist even for large $T_{\delta}$, with $P_{w}$ localizing at $w=0$ and $w=+1$ in the defect time interval $T_{\delta}/T_{d}\in\left[1.2,1.5\right]$, and at $w=0$ and $w=-1$ for $T_{\delta}/T_{d}\in\left[1.7,2.0\right)$.
@@
To understand why these features appear for $\lambda_{0}>\lambda_{c}$, we study the dynamics of $S^{x}$ and the photon number, $|a|^{2}$ during the defect protocol. We present in Fig. 7(a) the dynamics of these two quantities for $N=5000$ when $T_{\delta}=T_{d}$, which is close to the discontinuous transition point in $T_{\delta}\approx 0.8T_{d}$. Notice that unlike the DTC states for $\lambda_{0}<\lambda_{c}$, $S^{x}$ does not go to zero throughout the defect. Instead, it retains a nonzero value, implying that as soon as the defect protocol is turned on, the system reverts to the SP. This is further highlighted in the bottom panel of Fig. 7(a), which depicts that the photon number during the defect is larger than the threshold $|a|^{2}=10$ around which we expect the effects of the quantum fluctuations to become relevant. This explains not only the robustness of the bit-flip operation for $T_{\delta}$ close to $T_{\delta,c}=0.8T_{d}$, but also why the discontinuous transition at this critical point survives even for small $N$. We demonstrate in Fig. 7(b) however that as soon as we consider larger $T_{\delta}$, the defect protocol can push the system into a light-induced NP (LINP), in which both $S^{x}$ and $|a|^{2}$ approaches zero [79, 80]. This results in the system becoming significantly affected by the quantum noise, reducing the overall robustness of the bit-flip operation.
@@
Next, to explain the splitting of the $P_{w}$ in Fig. 6, let us again consider the dynamics shown in Fig. 7(a). We note that depending on the state of the system before the defect protocol at $t=0$ for $T_{\delta}/T_{d}\in\left[0.8,1.2\right)$, $S^{x}$ can either oscillate around a positive or a negative value during the defect protocol, consistent with the $\mathbb{Z}_{2}$-symmetry breaking nature of the SP. Due to this additional degree of freedom, the complex amplitude of $S^{x}$ can revolve around $A_{R}=0$ either clockwise or counter-clockwise, leading to the splitting of the half-winding distribution to $w=+1$ and $w=-1$. For ranges of defect duration given by $T_{\delta}/T_{d}\in\left[1.2,1.5\right]$, and $T_{\delta}/T_{d}\in\left[1.7,2.0\right)$, both the $S^{x}$ and $|a|^{2}$ are close to zero. This leads to quantum fluctuations removing the initial memory of the system, and thus, at $t>T_{\delta}$, the fluctuations can either push the system to its original oscillation pattern or send it to a new DTC state. This then explains the splitting of the half-winding distribution at $w=0$ and $w=\pm 1$, depending on the value of $T_{\delta}$. Note that this bit-flip mechanism is different from the cases of the PD and DTC states for $\lambda_{0}<\lambda_{c}$, where the oscillation becomes macroscopic as soon as $T_{\delta}>T_{\delta,c}$, as shown for instance in Fig. 1(h), which then restricts $w$ to a single value.
@@
To summarize our results in this section, we demonstrate the robustness of bit-flip operations on PD and DTC states using the defect protocol in Eq. (14) even for strong dissipation and noise strength, which is of thermal nature for the DPP and quantum origin for the ODM. Moreover, we have also shown that we can still switch DTC states even at parameter regimes where the system cannot be mapped onto a standard PO. In particular, depending on how the system responds during the defect protocol, its transient state may even protect it from quantum fluctuations, leading to a more robust bit-flip operation. Due to the possibility of quenching the system into a new dynamical phase during the defect, we will now consider in the following section a generalization of the defect protocol that may allow us to fine-tune the system’s state during the defect to obtain a more precise bit-flip operation.
@@ IV Generalized defect protocol
Figure 8: (a) Phase diagram of the ODM for $\lambda_{0}=1.1\lambda_{c}$, and $\kappa=1.0\omega$ in the absence of any defects. (b)-(c) Half-winding number distribution of the ODM under the quenched-frequency protocol for (b) $\omega_{d}^{\prime}=0.9\omega$, (c) $\omega_{d}^{\prime}=1.2\omega$, and (d) $\omega_{d}^{\prime}=1.45\omega$. The dark circles in (a) marks the initial choice of $\omega_{d}=0.8\omega$ and the remaining markers denote the driving frequencies during the defect protocol $\omega_{d}^{\prime}$ used in (b)-(d). The remaining driving parameters are $A=0.55$, and $N=5000$.
As discussed in Sec. III.3, by applying the defect protocol according to Eq. (14) on a DTC state for $\lambda_{0}>\lambda_{c}$, the ODM can be pushed into a new phase while the defect is switched on. To explain this behavior, note that Eq. (14) can also be viewed as a sudden quench protocol of the driving frequency from $\omega_{d}$ to $\omega_{d}^{\prime}=\omega_{d}+2\pi/T_{\delta}$. It is then fixed at $\omega^{\prime}$ for some duration $T_{\delta}$ until the driving frequency is reverted to $\omega_{d}$ at $t>T_{\delta}$. In Fig. 8(a), we show that the dynamical phase of the ODM for $\lambda_{0}>\lambda_{c}$ depends on $\omega_{d}$. As such, applying a defect effectively pushes the system onto a new phase if $T_{\delta}$ is sufficiently large for the system to relax to the stable phase at a given $\omega_{d}^{\prime}$. In the following, we exploit this driving-controlled quench of the dynamical phase to switch the bit encoded in the DTC states.
We now consider a modification of the defect protocol, which now corresponds to quenching the driving frequency
@@
where we set $\omega_{d}^{\prime}=\omega_{d}+2\pi/T_{\delta}$ to be a constant value, while we treat the duration of the defect $T_{r}$ as a separate tuning parameter. Unlike in Eq. (14), the protocol in Eq. (21) is only continuous for $T_{r}=nT_{\delta}$, with $n\in\mathbb{Z}$, otherwise the phase of the drive jumps from $2\pi T_{r}/T_{\delta}$ to zero after the defect protocol. We consider a continuous version of this protocol in Appendix C.
@@
We present in Figs. 8(b)-8(d) the $P_{w}$ of the ODM for $\lambda_{0}>\lambda_{c}$ and varying $\omega_{d}^{\prime}$ representing quenches into three different states: a quench to a DTC with a different response frequency; a quench to the LINP; and a quench to the SP, as shown in Figs. 8(b)-8(d), respectively. We observe that among the three scenarios, we see a more localized distribution when we quench our system to another DTC phase with a different response frequency. In this case, the $P_{w}$ alternates between even and odd $w$, forming a staircase localized only at integer values of $w$. This behavior implies that the success of the bit-flip operation for this protocol only depends on $T_{r}$, making it a timing-based method for performing bit-flips.
@@
For quenches into the LINP, as exemplified in Fig. 8(c), we observe a spreading of the half-winding distribution, which becomes more prominent as we increase $T_{r}$. This can be attributed to the thermalization of the system as it relaxes into the LINP during the defect protocol. As a result, the success of the bit-flip operation will highly depend on its state after $t=T_{r}$, which is then set by the quantum fluctuations present at that particular time. Note that while the $P_{w}$ spreads for the LINP case, it remains localized in integer values of $w$ due to the particular form of the protocol in Eq. (21), in which the driving phase reverts to zero at $t>T_{r}$. This restricts the dynamics after the defect protocol to the original degenerate DTC states, thereby leading to a difference between the initial and final absolute phase of $\triangle\varphi\equiv\varphi(t\rightarrow\infty)-\varphi(0)=\pi w$, with $w\in\mathbb{Z}$.
@@
Finally, for the case of the quench into the SP, we show in Fig. 8(d) that the $P_{w}$ spreads similarly to that for the LINP case. We can attribute this behavior to our particular choice of $\omega_{d}^{\prime}$, wherein as shown in Fig. 8(a), our chosen $\omega_{d}^{\prime}$ is close to the critical line separating the SP and the LINP. As a result, the photon number during the defect protocol is not macroscopic enough to overcome quantum fluctuations, leading to a delocalized $P_{w}$.
@@
Figure 9: (a)-(b) Half-winding number distribution of the ODM under the switch-off protocol for the driving parameters (a) $\{\lambda_{0},\omega_{d},A\}=\{0.9\lambda_{c},2\omega_{r,\mathrm{ODM}},A_{r,\mathrm{ODM}}+0.1\}$ and (b) $\{\lambda_{0},\omega_{d},A\}=\{1.1\lambda_{c},0.8\omega,0.55\}$. The cavity dissipation for both figures is $\kappa=1.0\omega$.
@@
We can further explore the robustness of the bit-flip operation via quenches towards the SP by considering the case of $\omega_{d}^{\prime}\rightarrow 0$, where the system is pushed back to the corresponding equilibrium phase prior to the driving. We present in Figs. 9(a) and 9(b) the $P_{w}$ when $\omega_{d}^{\prime}=0$ for $\lambda_{0}=0.9\lambda_{c}$, and $\lambda_{0}=1.1\lambda_{c}$, respectively. We demonstrate that for quenches back to the NP, $\lambda_{0}<\lambda_{c}$, the $P_{w}$ spreads out as $T_{r}$ increases, consistent with our results for quenches into the LINP. On the other hand, for $\lambda_{0}>\lambda_{c}$, the $P_{w}$ finally localizes to either an even or odd integer of $w$, signaling that robust bit-flip operations can be performed by a sudden quench into the SP. In this case, the parameters considered lead to a steady state with macroscopic excitation of the cavity mode enough to suppress the quantum fluctuations.
@@
Our results here demonstrate that bit-flip operations can be implemented by quenching the system from a DTC to another phase and then bringing the system back to a DTC after some time has elapsed. This bit-flip operation becomes even more robust when the system is quenched into a phase that has enough macroscopic excitations to counteract the effects of the quantum noise. Note that this mechanism also explains the robustness of the bit-flip operations for both the PD and DTC states using the protocol in Eq. (14) for $T_{\delta}\gg T_{d}$.
@@ V Summary and Discussion
In this work, we have proposed that a dynamical protocol based on introducing a defect in the phase of the drive [8] can be utilized for bit-flip operations of classical bits encoded in the PD and DTC states of open systems. Furthermore, we have demonstrated the robustness of the bit-flip protocol against thermal and quantum noise.
@@
We primarily focused on two systems: (i) a classical parametrically driven pendulum connected to a thermal bath and; (ii) a quantum system of two-level systems or qubits coupled to a single photonic mode described by the open Dicke model. The ODM allows for investigating two different phases dependent on the strength of the light-matter interaction. For $\lambda_{0}<\lambda_{c}$, the steady state of the ODM is the NP, which can be approximated as a coupled PO when periodically driven. Meanwhile, for $\lambda_{0}>\lambda_{c}$, the ODM spontaneously breaks its $\mathbb{Z}_{2}$ symmetry in the static limit, leading to the breakdown of its coupled POs picture. The latter case allows us to explore bit-flip operations beyond the framework of parametric oscillators.
@@
We have demonstrated that in the noiseless limit, which corresponds to the zero-temperature limit ($T\rightarrow 0$) for the DPP and the thermodynamic limit ($N\rightarrow\infty$) for the ODM with $\lambda_{0}<\lambda_{c}$, the bit-flip operations become robust with increasing dissipation strength. In particular, for weak dissipation, their respective bit-flip diagrams exhibit fractal structures that can be attributed to the two systems entering transient irregular dynamics after applying the defect protocol. These fractal structures vanish for strong dissipation, allowing for robust switching in a large area of the parameter space spanned by $T_{\delta}$ and $A$. In the presence of fluctuations, the bit-flip operation remains robust against fluctuations when $T_{\delta}$ is sufficiently far from the critical defect duration needed for a successful bit-flip. Also, for $T_{\delta}\lessapprox T_{\delta,c}$, the half-winding number reveals that the defect protocol can have a non-zero probability to switch the PD and DTC states, leading to a crossover-like behavior of the switching probability near $T_{\delta,c}$. This is in stark contrast with the discontinuous transition observed in noiseless POs [8].
@@
For the case of $\lambda_{0}>\lambda_{c}$ in the ODM, we have shown that the system can be pushed into various phases when the defect is switched on for a sufficient time. This has motivated us to explore the possibility of implementing bit-flip operations by quenching the driving frequency with variable defect duration $T_{r}$. We have shown that robust bit-flip operations are still possible if we choose an appropriate $\omega_{d}^{\prime}$ that quenches the system into a new phase with a macroscopic occupation number enough to suppress the effects of noise from the dissipative channel. This mechanism is also responsible for protecting the bit-flip operation for large $T_{\delta}$ for both the thermal DPP and the ODM for $\lambda_{0}<\lambda_{c}$.
@@
Our work provides a general framework for implementing bit-flip operations on PD and DTC states on systems where thermal and quantum fluctuations become relevant to the system’s dynamics. We have extended the notion of defect protocols based on ramping the phase of the drive and have addressed subtle points concerning the specific form of the protocol for robust bit-flips. Concomitantly, we have provided a way to dynamically manipulate DTCs, which could be useful for more in-depth exploration on the nature of DTCs and potential applications of PD states and DTCs as logical bits.

### v2

@@ III.3 Beyond parametric oscillator models
Figure 5: (a) Switching probability of the ODM for different particle number and for the parameters $\lambda_{0}=1.1\lambda_{c}$, $\kappa=1.0\omega$, $\omega_{d}=0.8\omega$, and $A=0.55$. (b)-(d) Half-winding number distribution of the ODM for (b) $N=10^{4}$, (c) $N=10^{5}$, (d) $N=10^{6}$. The same set of parameters in (a) is used to construct the histograms in (b)–(e).
We present in Fig. 5(a) the bit-flip probability of the DTC states in the $\lambda_{0}>\lambda_{c}$ regime. Notice that the $P_{s}$ has a more complex behavior, with its most striking feature being the discontinuity at $T_{\delta,c}\approx 0.8T_{d}$ that persists even for small $N$. This is in contrast to the behavior observed so far with the thermal PO and DTC states in the NP in Figs. 3(a) and 3(b), respectively. As we further increase $T_{\delta}$, however, $P_{s}$ drops to a lower value. In particular, when we neglect the temporal fluctuations and only consider the initial noise, $P_{s}\approx 0.5$ within the range of $T_{\delta}/T_{d}\in\left[1.2,1.5\right]$ and $T_{\delta}/T_{d}\in\left[1.7,2.0\right)$, while it goes to zero for $T_{\delta}/T_{d}\in\left(1.5,1.7\right)$. Unlike the discontinuous transition in $T_{\delta}=0.8T_{d}$ however, this feature does not persist for small $N$, and instead becomes smoothened out similar to the crossovers observed in Figs. 3(a) and 3(b).
From the distribution of $w$ shown in Figs. 5(b)–5(d), we see that at $T_{\delta,c}=0.8T_{d}$, the distribution jumps from $w=0$ to $+1$ and then splits into $w=+1$ and $-1$ as we further increase $T_{\delta}$. This splitting of the $P_{w}$ persists even for large values of $N$, which, again, is in contrast to the thermal PO and the DTC states for $\lambda_{0}<\lambda_{c}$, where the distribution only becomes localized at one value of $w$. This splitting persists even for large $T_{\delta}$, with $P_{w}$ localizing at $w=0$ and $+1$ in the defect time interval $T_{\delta}/T_{d}\in\left[1.2,1.5\right]$, and at $w=0$ and $-1$ for $T_{\delta}/T_{d}\in\left[1.7,2.0\right)$.
@@
To understand why these features appear for $\lambda_{0}>\lambda_{c}$, we study the dynamics of $S^{x}$ and the photon number, $|a|^{2}$ during the defect protocol. We present in Fig. 6(a) the dynamics of these two quantities for $N=5000$ when $T_{\delta}=T_{d}$, which is close to the discontinuous transition point in $T_{\delta}\approx 0.8T_{d}$. Notice that unlike the DTC states for $\lambda_{0}<\lambda_{c}$, $S^{x}$ does not go to zero throughout the defect. Instead, it retains a nonzero value, implying that as soon as the defect protocol is turned on, the system reverts to the SP. This is further highlighted in the bottom panel of Fig. 6(a), which depicts that the photon number during the defect is larger than the threshold $|a|^{2}=10$ around which we expect the effects of the quantum fluctuations to become relevant. This explains not only the robustness of the bit-flip operation for $T_{\delta}$ close to $T_{\delta,c}=0.8T_{d}$, but also why the discontinuous transition at this critical point survives even for small $N$. We demonstrate in Fig. 6(b), however, that as soon as we consider larger $T_{\delta}$, the defect protocol can push the system into a light-induced NP, in which both $S^{x}$ and $|a|^{2}$ approach zero [69, 70]. This results in the system becoming significantly affected by the quantum noise, reducing the overall robustness of the bit-flip operation.
@@
Next, to explain the splitting of the $P_{w}$ in Fig. 5, let us again consider the dynamics shown in Fig. 6(a). We note that depending on the state of the system before the defect protocol at $t=0$ for $T_{\delta}/T_{d}\in\left[0.8,1.2\right)$, $S^{x}$ can either oscillate around a positive or a negative value during the defect protocol, consistent with the $\mathbb{Z}_{2}$-symmetry breaking nature of the SP. Due to this additional degree of freedom, the complex amplitude of $S^{x}$ can revolve around $A_{R}=0$ either clockwise or counter-clockwise, leading to the splitting of the half-winding distribution to $w=+1$ and $-1$. For ranges of defect duration given by $T_{\delta}/T_{d}\in\left[1.2,1.5\right]$, and $T_{\delta}/T_{d}\in\left[1.7,2.0\right)$, both the $S^{x}$ and $|a|^{2}$ are close to zero. This leads to quantum fluctuations removing the initial memory of the system, and thus, at $t>T_{\delta}$, the fluctuations can either push the system to its original oscillation pattern or send it to a new DTC state. This then explains the splitting of the half-winding distribution at $w=0$ and ±$\pm 1$, depending on the value of $T_{\delta}$. Note that this bit-flip mechanism is different from the cases of the period-doubled states and DTC states for $\lambda_{0}<\lambda_{c}$, where the oscillation becomes macroscopic as soon as $T_{\delta}>T_{\delta,c}$, as shown for instance in Fig. 1(h), which then restricts $w$ to a single value.
@@
To summarize our results in this section, we demonstrate the robustness of bit-flip operations on period-doubled and DTC states using the defect protocol in Eq. (12) even for strong dissipation and noise strength, which is of thermal nature for the PO and quantum origin for the ODM. Moreover, we have also shown that we can still switch DTC states even at parameter regimes where the system cannot be mapped onto a standard PO. In particular, depending on how the system responds during the defect protocol, its transient state may even protect it from quantum fluctuations, leading to a more robust bit-flip operation. Due to the possibility of quenching the system into a new dynamical phase during the defect, we will now consider in the following section a generalization of the defect protocol that may allow us to fine-tune the system’s state during the defect to obtain a more precise bit-flip operation.
@@ IV Generalized defect protocol
Figure 7: (a) Phase diagram of the ODM for $\lambda_{0}=1.1\lambda_{c}$, and $\kappa=1.0\omega$ in the absence of any defects. (b), (c) Half-winding number distribution of the ODM under the quenched-frequency protocol for (b) $\omega_{d}^{\prime}=0.9\omega$, (c) $\omega_{d}^{\prime}=1.2\omega$, and (d) $\omega_{d}^{\prime}=1.45\omega$. The dark circles in (a) marks the initial choice of $\omega_{d}=0.8\omega$ and the remaining markers denote the driving frequencies during the defect protocol $\omega_{d}^{\prime}$ used in (b)–(d). The remaining driving parameters are $A=0.55$, and $N=5000$.
As discussed in Sec. III.3, by applying the defect protocol according to Eq. (12) on a DTC state for $\lambda_{0}>\lambda_{c}$, the ODM can be pushed into a new phase while the defect is switched on. To explain this behavior, note that Eq. (12) can also be viewed as a sudden quench of the driving frequency from $\omega_{d}$ to $\omega_{d}^{\prime}=\omega_{d}+2\pi/T_{\delta}$. It is then fixed at $\omega_{d}^{\prime}$ for some duration $T_{\delta}$ until the driving frequency is reverted to $\omega_{d}$ at $t>T_{\delta}$. In Fig. 7(a), we show that the dynamical phase of the ODM for $\lambda_{0}>\lambda_{c}$ depends on $\omega_{d}$. As such, applying a defect effectively pushes the system onto a new phase if $T_{\delta}$ is sufficiently large for the system to relax to the stable phase at a given $\omega_{d}^{\prime}$. In the following, we exploit this driving-controlled quench of the dynamical phase to switch the bit encoded in the DTC states.
We now consider a modification of the defect protocol, which corresponds to quenching the driving frequency
@@
where we set $\omega_{d}^{\prime}=\omega_{d}+2\pi/T_{\delta}$ to be a constant value, while we treat the duration of the defect $T_{r}$ as a separate tuning parameter. Unlike in Eq. (12), the protocol in Eq. (19) is only continuous for $T_{r}=nT_{\delta}$, with $n\in\mathbb{Z}$, otherwise the phase of the drive jumps from $2\pi T_{r}/T_{\delta}$ to zero after the defect protocol. We consider a continuous version of this protocol in Appendix D.
@@
We present in Figs. 7(b)–7(d) the $P_{w}$ of the ODM for $\lambda_{0}>\lambda_{c}$ and varying $\omega_{d}^{\prime}$ representing quenches into three different states: a quench to a DTC with a different response frequency; a quench to the light-induced NP; and a quench to the SP, as shown in Figs. 7(b)–7(d), respectively. We observe that among the three scenarios, we see a more localized distribution when we quench our system to another DTC phase with a different response frequency. In this case, the $P_{w}$ alternates between even and odd $w$, forming a staircase localized only at integer values of $w$. This behavior implies that the success of the bit-flip operation for this protocol only depends on $T_{r}$, making it a timing-based method for performing bit-flips.
@@
For quenches into the light-induced NP, as exemplified in Fig. 7(c), we observe a spreading of the half-winding distribution, which becomes more prominent as we increase $T_{r}$. This can be attributed to the thermalization of the system as it relaxes into the light-induced NP during the defect protocol. As a result, the success of the bit-flip operation will highly depend on its state after $t=T_{r}$, which is then set by the quantum fluctuations present at that particular time. Note that while the $P_{w}$ spreads for the light-induced NP case, it remains localized in integer values of $w$ due to the particular form of the protocol in Eq. (19), in which the driving phase reverts to zero at $t>T_{r}$. This restricts the dynamics after the defect protocol to the original degenerate DTC states, thereby leading to a difference between the initial and final absolute phase of $\triangle\varphi\equiv\varphi(t\rightarrow\infty)-\varphi(0)=\pi w$, with $w\in\mathbb{Z}$.
@@
Finally, for the case of the quench into the SP, we show in Fig. 7(d) that the $P_{w}$ spreads similarly to that for the light-induced NP case. We can attribute this behavior to our particular choice of $\omega_{d}^{\prime}$, wherein as shown in Fig. 7(a), our chosen $\omega_{d}^{\prime}$ is close to the critical line separating the SP and the light-induced NP. As a result, the photon number during the defect protocol is not macroscopic enough to overcome quantum fluctuations, leading to a delocalized $P_{w}$.
@@
Figure 8: (a), (b) Half-winding number distribution of the ODM under the switch-off protocol for the driving parameters (a) $\{\lambda_{0},\omega_{d},A\}=\{0.9\lambda_{c},2\omega_{r,\mathrm{ODM}},A_{r,\mathrm{ODM}}+0.1\}$ and (b) $\{\lambda_{0},\omega_{d},A\}=\{1.1\lambda_{c},0.8\omega,0.55\}$. The cavity dissipation for both figures is $\kappa=1.0\omega$.
@@
We can further explore the robustness of the bit-flip operation via quenches towards the SP by considering the case of $\omega_{d}^{\prime}\rightarrow 0$, where the system is pushed back to the corresponding equilibrium phase prior to the driving. We present in Figs. 8(a) and 8(b) the $P_{w}$ when $\omega_{d}^{\prime}=0$ for $\lambda_{0}=0.9\lambda_{c}$, and $\lambda_{0}=1.1\lambda_{c}$, respectively. We demonstrate that for quenches back to the NP, $\lambda_{0}<\lambda_{c}$, the $P_{w}$ spreads out as $T_{r}$ increases, consistent with our results for quenches into the light-induced NP. On the other hand, for $\lambda_{0}>\lambda_{c}$, the $P_{w}$ finally localizes to either an even or odd integer of $w$, signaling that robust bit-flip operations can be performed by a sudden quench into the SP. In this case, the parameters considered lead to a steady state with macroscopic excitation of the cavity mode enough to suppress the quantum fluctuations.
@@
Our results here demonstrate that bit-flip operations can be implemented by quenching the system from a DTC to another phase and then bringing the system back to a DTC after some time has elapsed. This bit-flip operation becomes even more robust when the system is quenched into a phase that has enough macroscopic excitations to counteract the effects of the quantum noise. Note that this mechanism also explains the robustness of the bit-flip operations for both the period-doubled and DTC states using the protocol in Eq. (12) for $T_{\delta}\gg T_{d}$.
@@ V Summary and Discussion
In this work, we have proposed that a protocol based on introducing a defect in the phase of the drive [8] can be utilized for bit-flip operations of classical bits encoded in the period-doubled and DTC states of open systems. Furthermore, we have demonstrated the robustness of the bit-flip protocol against thermal and quantum noise.
@@
We primarily focused on two systems: (i) a classical parametric oscillator connected to a thermal bath and; (ii) a quantum system of two-level systems or qubits coupled to a single photonic mode described by the open Dicke model. The ODM allows for investigating two different phases dependent on the strength of the light-matter interaction. For $\lambda_{0}<\lambda_{c}$, the steady state of the ODM is the NP, which can be approximated as a coupled PO when periodically driven. Meanwhile, for $\lambda_{0}>\lambda_{c}$, the ODM spontaneously breaks its $\mathbb{Z}_{2}$ symmetry in the static limit, leading to the breakdown of its coupled POs picture. The latter case allows us to explore bit-flip operations beyond the framework of parametric oscillators.
@@
We have demonstrated that in the noiseless limit, which corresponds to the zero-temperature limit ($T\rightarrow 0$) for the PO and the thermodynamic limit ($N\rightarrow\infty$) for the ODM with $\lambda_{0}<\lambda_{c}$, the bit-flip operations become robust with increasing dissipation strength. In particular, for weak dissipation, their respective bit-flip diagrams exhibit fractal structures that can be attributed to the two systems entering transient irregular dynamics after applying the defect protocol. These fractal structures vanish for strong dissipation, allowing for robust switching in a large area of the parameter space spanned by $T_{\delta}$ and $A$. In the presence of fluctuations, the bit-flip operation remains robust against fluctuations when $T_{\delta}$ is sufficiently far from the critical defect duration needed for a successful bit-flip. Also, for $T_{\delta}\lessapprox T_{\delta,c}$, the half-winding number reveals that the defect protocol can have a non-zero probability to switch the period-doubled and DTC states, leading to a crossover-like behavior of the switching probability near $T_{\delta,c}$. This is in stark contrast with the discontinuous transition observed in noiseless POs [8].
@@
For the case of $\lambda_{0}>\lambda_{c}$ in the ODM, we have shown that the system can be pushed into various phases when the defect is switched on for a sufficient time. This has motivated us to explore the possibility of implementing bit-flip operations by quenching the driving frequency with variable defect duration $T_{r}$. We have shown that robust bit-flip operations are still possible if we choose an appropriate $\omega_{d}^{\prime}$ that quenches the system into a new phase with a macroscopic occupation number enough to suppress the effects of noise from the dissipative channel. This mechanism is also responsible for protecting the bit-flip operation for large $T_{\delta}$ for both the thermal PO and the ODM for $\lambda_{0}<\lambda_{c}$.
@@
Our work provides a general framework for implementing bit-flip operations on period-doubled and DTC states on systems where thermal and quantum fluctuations become relevant to the system’s dynamics. We have extended the notion of defect protocols based on ramping the phase of the drive and have addressed subtle points concerning the specific form of the protocol for robust bit-flips. Concomitantly, we have provided a general method to dynamically manipulate DTCs, which can be of interest for more in-depth exploration of their nature, and can be used as a standard tool for potential applications of DTCs such as, but not limited to, quantum metrology [71, 72, 73] and quantum computing [74]. Our results can be tested in existing platforms emulating the parametric oscillators [2, 9, 7], and also systems that can emulate the open Dicke model, such as the atom-cavity system [44, 75, 76], ensembles of nitrogen-vacancy centers [77, 78], and cavity-magnon systems [79], to name a few.
A natural extension of our work is to explore the effects of finite temperature in the robustness of the bit-flip operation on the DTC states. Also, given the generality of the protocol, it will be interesting to test this in systems that can emulate higher-order DTCs, in which the period of the DTC is greater than twice of the driving period, resulting to multiple degenerate DTC states. Higher-order DTCs have been explored in bosonic systems [80, 51] and spin systems such as the Lipkin-Meshkov-Glick model [30], and its generalization the $p$-spin model [16, 81]. This dynamical phase has also been observed experimentally in Rydberg atoms [32] and Kerr resonators [82]. It will also be interesting to know the physical interpretation behind the half-winding number associated with the bit-flips, and whether one can relate it to topological phenomena that may arise in higher-order DTCs, in spirit of the topological interpretation of $w$ in Refs. [8, 54].